ATA UN_HIBERNATION - REPORT 4





I am here on my fourth trip to the ATA to help with un-hibernation efforts. The plan is to have our SETI searches operating by the end of the month. We will be announcing the exact date some time next week. At least that is the current plan.

The temperatures have warmed up a bit. This is good - several days ago it was raining. Today we have mostly clear skys. 

Mount lassen seemed to have a bit more snow than the last time. Mount Shasta was barely visible due to a purplish haze.

 

 

 

 

 

 

 

 

Mount Lassen (left), Mount Shasta (right)

We are in the process of hooking up the computer systems. This involves connecting up 14 servers donated to the project by Dell.

The Dell 6100 Servers

All these servers produce a lot of heat. So much heat that the Signal Processing Room (SPR) that houses all the computer systems will quickly overheat. Work is being done to increase airflow. The first of this work is to mount special thick plastic curtains in the SPR that will guide airflow to provide better cooling to certain areas.

In the attic - installing ceiling supports for the curtains

One thing that comes with all this work is a mess.

The Control Room 2 weeks ago (left), and today (right)

And on top of all, we had a school from Susanville, CA tour the site. A busy day and it is only 10am!